### v1

# Privacy-protected Retrieval-Augmented Generation for Knowledge Graph Question Answering

Abstract
Large Language Models (LLMs) often suffer from hallucinations and outdated or incomplete knowledge. Retrieval-Augmented Generation (RAG) is proposed to address these issues by integrating external knowledge like that in knowledge graphs (KGs) into LLMs. However, leveraging private KGs in RAG systems poses significant privacy risks due to the black-box nature of LLMs and potential insecure data transmission, especially when using third-party LLM APIs lacking transparency and control. In this paper, we investigate the privacy-protected RAG scenario for the first time, where entities in KGs are anonymous for LLMs, thus preventing them from accessing entity semantics. Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless identifiers of anonymous entities. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information. (2) How to retrieve question-relevant anonymous entities.
To address these challenges, we propose a novel Abstraction Reasoning on Graph (ARoG) framework including relation-centric abstraction and structure-oriented abstraction strategies. For challenge (1), the first strategy abstracts entities into high-level concepts by dynamically capturing the semantics of their adjacent relations. Hence, it supplements meaningful semantics which can further support the retrieval process. For challenge (2), the second strategy transforms unstructured natural language questions into structured abstract concept paths. These paths can be more effectively aligned with the abstracted concepts in KGs, thereby improving retrieval performance. In addition to guiding LLMs to effectively retrieve knowledge from KGs, these two abstraction strategies also strictly protect privacy from being exposed to LLMs. Experiments on three datasets demonstrate that ARoG achieves strong performance and privacy-robustness, establishing a new practical direction for privacy-protected RAG systems.
1 Introduction
Figure 1: Comparison between previous RAG (Step A) with privacy-protected RAG (Step I-II) system for KGQA.
Large Language Models (LLMs) exhibit remarkable capabilities across a wide range of natural language tasks (Wang et al., 2024). However, their practical deployment remains constrained by critical limitations, including factual inaccuracies (hallucinations) and outdated or incomplete knowledge (Ji et al., 2023). To address these issues, the Retrieval-Augmented Generation (RAG) technique has been proposed to first retrieve factual knowledge from external sources and then incorporate them into LLMs during inference, thereby enhancing the factual grounding and reliability of LLM outputs (Shi et al., 2024).
KGs can store a large amount of factual knowledge. Hence, they are widely used in the RAG system as external knowledge sources to provide supplementary knowledge to LLMs (Sun et al., 2024). The effectiveness of KG-enhanced RAG systems is commonly evaluated through Knowledge Graph Question Answering (KGQA) task. Specifically, given a question, the RAG system aims first to retrieve relevant factual triplets from the KGs and then generate answers based on them.
In practice, many private KGs contain sensitive information like personal data and confidential company details (Chen et al., 2023). When using such private KGs to answer questions requiring privacy details, current RAG systems must retrieve relevant privacy data and expose them to the LLMs. As shown in the Step A of Figure 1, when retrieving triplets to answer the question “Where does Bronny live?”, the factual triplet (Bronny,livesin,L.A.)(Bronny,lives\ in,L.A.) containing sensitive information will be exposed to LLMs. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their private data are collected, stored, or processed.
In this paper, we make the first attempt to address privacy concerns when the RAG system retrieves external knowledge from KGs. In this privacy-protected RAG scenario, as depicted in the Step I of Figure 1, entities within the KG are anonymous for LLMs and replaced with corresponding unique Machine Identifiers (MIDs), which are encrypted and contain no semantic content. As a result, as illustrated in the Step II of Figure 1, the LLMs are unable to access the semantic information (e.g., the specific types, names, and descriptions) of these entities.
Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless MIDs. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information. (2) How to retrieve question-relevant anonymous entities.
To address these challenges, we propose a novel Abstraction Reasoning on Graph (ARoG) framework including relation-centric abstraction and structure-oriented abstraction strategies.
For challenge (1), the relation-centric abstraction strategy treats the anonymous entities as subject or object nouns and their adjacent relations as predicate verbs, then abstracts the entities into corresponding higher-level concepts. These concepts are generated by LLMs based on the adjacent relations of entities. This strategy captures their in-context semantics and then appending to the MIDs, thereby overcoming challenge (1). For instance, an entity that serves as the subject of relations “time_zones”, “contained_by”, “population” and the object of relation “citytown” can be abstracted into “geographic location”.
For challenge (2), the structure-oriented abstraction strategy transforms unstructured natural language questions into structured abstract concept paths. These paths can be more effectively aligned with the abstracted concepts in KGs than naive questions, thereby improving retrieval performance. For instance, the abstract concept path to the question “What is the name of the daughter of the artist who had the The Mrs. Carter Show World Tour?” is “Nicki Minaj (artist) $\rightarrow$ had $\rightarrow$ The Mrs. Carter Show World Tour; Nicki Minaj (artist) $\rightarrow$ has daughter named $\rightarrow$ Chiara Fattorini (person)”. Notably, the semantic similarity between this path and relevant triplets persists even with inaccurate entities, as the path captures the relational structure of abstracted concepts.
To fulfill the complete privacy-protected RAG, our ARoG framework also incorporates Abstraction-driven Retrieval module and Generator module. These two modules ensure robust performance in retrieving question-relevant triplets from KGs and generating answers. To evaluate the effectiveness of our ARoG framework, we conduct experiments in the privacy-protected RAG scenario on three popular yet diverse datasets including WebQSP, CWQ, and GrailQA. Our ARoG framework achieves the state-of-the-art (SoTA) performance on all three datasets.
Our contributions can be summarized as follows.
  • We explore the privacy-protected RAG scenario for the first time, which aims to address privacy concerns when the RAG system retrieves external knowledge from KGs.
  • We propose a novel RAG-based framework ARoG, utilizing two abstraction strategies to tackle challenges in the privacy-protected RAG scenario while balancing high performance and data privacy.
  • We conduct extensive experiments across three public datasets, providing empirical evidence of the effectiveness and robustness of our proposed ARoG.
2 Related Work
In this section, we review related works for the KGQA task.
2.1 Semantic Parsing for KGQA
Semantic parsing (SP)-based methods treat LLMs as semantic parsers that translate questions into formal queries using labeled exemplars. To address errors in the initial formal queries, Li et al. (2023) utilize entity and relation binders grounded in KGs, Nie et al. (2024) reformulate the generation process into code generation process; and Zhang et al. (2025) adapt generation process to function calling using Condition Graphs (CGs), a variant of KGs. However,these methods do not fundamentally resolve the challenges posed by formatting errors, and the performance is heavily limited by the quality and quantity of labeled exemplars.
2.2 Retrieval-Agumented Generation for KGQA
Compared with SP-based methods, RAG-based methods usually have superior performance. RAG-based methods employ LLMs as retrievers to fetch question-relevant triplets as evidence, and employ LLMs as generators to infer answers with evidence. To improve the quality of evidence, Jiang et al. (2023a) propose an iterative LLM-based “Read-then-Reason” framework; Sun et al. (2024) utilize LLMs to perform beam search on KGs; and Xu et al. (2024) utilize LLMs to supplement additional knowledge. However, previous RAG-based KGQA methods directly feed factual triplets within KGs to LLMs, which may undermine data privacy. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their privacy data are collected, stored, or processed.
In this paper, we make the first attempt to explore the privacy-protected RAG scenario for KGQA, which aims to address privacy concerns when the RAG system retrieves external knowledge from KGs.
Figure 2: An overview of ARoG framework, which includes four primarily modules: Relation-centric Abstraction, Structure-oriented Abstraction, Abstraction-driven Retrieval and Generator modules. The Relation-centric Abstraction module consists of three steps: relation retrieval, relation filtering and entity abstraction.
3 Methodology
3.1 Preliminary
Knowledge Graphs (KGs)
Knowledge Graphs (KGs) contain abundant factual triplets (${e_{h}},r,e_{t}$), that is, $KG=\{(e_{h},r,e_{t})\mid e_{h},e_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $e_{h}$, $r$, and $e_{t}$ represent the head entity, the relation, and the tail entity, respectively. $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively.
Knowledge Graph Question Answering (KGQA)
Given an nature language question $q$, the topic entities $E_{q}\ =\ \left\{e_{q}^{1},e_{q}^{2},\cdot\cdot\cdot,e_{q}^{n}\right\}$ (the main entities asked in the question (Qiu et al., 2018)), and a knowledge graph $KG$, the task of KGQA is to predict the answer entities $A_{q}\ =\ \left\{e_{a}^{1},e_{a}^{2},\cdot\cdot\cdot,e_{a}^{n}\right\}$, which can answer the question. Following previous works (Sun et al., 2024), we assume that the topic entities and answer entities are labeled and linked to the corresponding entities in $KG$, i.e., $E_{q},A_{q}\subseteq\mathcal{E}$.
Privacy-protected RAG scenario for KGQA
To prevent the LLM from accessing sensitive information within the KG, we anonymize all entities by representing them as MIDs. Specifically, we leverage the Freebase database system (Bollacker et al., 2008) to map each entity to a unique, encrypted MID. These MIDs are meaningless to the LLM. Crucially, to maintain task feasibility, the names of the topic entities explicitly mentioned in the question are available.
3.2 Overview
In order to solve the challenges in the privacy-protected RAG scenario, we propose a novel Abstraction Reasoning on Graph (ARoG) framework. The architecture of ARoG, illustrated in Figure 2, follows a Retrieval-then-Generation pipeline. It comprises four primary modules: the first three constitute the retrieval phase, while the latter one belongs to the generation phase, as detailed below.
(1) Relation-centric Abstraction (RA): To overcome the challenge of How can anonymous entities be converted into retrievable information, we develop the Relation-centric Abstraction module. It utilizes LLMs to transform anonymous entities into abstract concepts and then appends to the anonymous identifies to supplement semantic.
(2) Structure-oriented Abstraction (SA): To overcome the challenge of How to retrieve question-relevant anonymous entities, we develop the Structure-oriented Abstraction module. It uses LLMs to convert unstructured questions into structured abstract concept paths to guide the retrieval.
(3) Abstraction-driven Retrieval (AR): To retrieve the question-relevant triplets from a large-scale KG, this module sequentially performs searching and pruning operations over the KG, driven by the two aforementioned abstraction modules.
(4) Generator: This module employs an LLM to infer answers based on the question and the question-relevant triplets derived from Abstraction-driven Retrieval.
3.3 Relation-centric Abstraction Module
The Relation-centric Abstraction module treats relations adjacent to entities as verbs and the entities themselves as subject or object nouns. Leveraging the generative capabilities of LLMs, we infer abstract concepts for the entities from these verbs, thereby enriching their semantics.
As illustrated in Part (a) of Figure 2, the Relation-centric Abstraction module operates on a question $q$, $n$ topic entities $E_{q}$ and the KG, and comprises three steps as follows.
(1) Relation Retrieval
First, starting from a set of $n$ topic entities, we explore the KG to extract $n$ sets of adjacent relations. For each relation cluster $R_{t}$, we utilize an LLM to identify the $W$ most question-relevant relations, which are denoted as $R_{t,opt}$. The process is expressed as follows.
$R_{\text{t,opt}}=\mathop{\arg\max}_{\begin{subarray}{c}R_{\text{t,opt}}\subseteq R_{t}\\ |R_{\text{t,opt}}|\leq W\end{subarray}}\text{LLM}(R_{t},q,Inst_{rr},E_{rr}),$ (1)
where $W$ is a hyperparameter representing the width of the retrieval, $Inst_{rr}$ describes the details of the relation retrieval task, $E_{rr}$ denotes an exemplar, and LLM denotes LLM call.
After retrieval, we gather $n\times W$ entity-relation pairs. Subsequently, for each pair, we identify its adjacent entity cluster, resulting in a total of $n\times W$ candidate entity clusters.
(2) Relation Filtering
Second, while the entities in the candidate clusters are meaningless, their adjacent relations provide rich in-context semantic information. For a representative entity $e$ in the candidate entity clusters, we treat its adjacent relations as verbs, denoted as $R_{v}$. To filter out irrelevant relations, we employ SentenceTransformer (Reimers and Gurevych, 2020) to select the top-$K$ most relevant relations, denoted as $R_{v,opt}$, based on their cosine similarity to the question $q$. The process is expressed as follows.
$R_{v,opt}=\mathop{\arg\max}_{\begin{subarray}{c}R_{v,opt}\subseteq R_{v}\\ |R_{v,opt}|\leq K\end{subarray}}\sum_{r\in R_{v,opt}}\cos(\text{Emb}(q),\text{Emb}(r)),$ (2)
where Emb(.) represents the embedding results calculated by SentenceTransformer, and $K$ is set to 5 empirically.
(3) Entity Abstraction
To abstract the representative anonymous entity $e$, we first prompt an LLM to infer its corresponding concept based on the related relations $R_{v,opt}$. We then append this abstract concept to the entity’s MID, yielding the abstracted entity $e^{abs}$. The procedure is expressed as follows.
$e^{abs}=\text{LLM}(R_{v,opt},e,Inst_{ea},E_{ea}),$ (3)
where $Inst_{ea}$ describes the details of the entity abstraction, and $E_{ea}$ denotes a single exemplar for guidance.
Since the relations within the KG define the schema structures rather than sensitive information, sharing these relations with LLMs poses minimal privacy risk. Additionally, to minimize LLM calls, we assume that entities within the same entity cluster share a common concept, in line with Jiang et al. (2023b). If an entity is associated with multiple generated abstract concepts, these concepts are combined into a comma-separated sequence.
This abstraction process is applied to each triple $t=(e_{h},r,e_{t})$ in the KG, replacing its entities with their abstracted counterparts to form an abstracted triple $t^{abs}=(e_{h}^{abs},r,e_{t}^{abs})$. We then apply this transformation uniformly to the entire set of triplets $T_{q}=\{t_{1},t_{2},\cdot\cdot\cdot\}$, yielding its abstracted counterpart, $T_{q}^{abs}=\{t_{1}^{abs},t_{2}^{abs},\cdot\cdot\cdot\}$.
3.4 Structure-oriented Abstraction Module
The Structure-oriented Abstraction module converts an unstructured question into a structured abstract concept path $P_{q}$. During this process, unknown entities are also abstracted into their corresponding concepts. Formally, $P_{q}$ comprises a set of generated triplets, denoted as $P_{q}=\{t_{p1},t_{p2},\cdot\cdot\cdot\}$. By guiding the retrieval of relevant triplets and entities via semantic matching, this path effectively isolates the LLM from the raw entities within the KG.
A key advantage of this strategy is that the effectiveness of abstract concept paths is not reliant on correct entities. For example, as shown in Part (b) of Figure 2, although the generated entities “Nicki Minaj” and “Chiara Fattorini” are incorrect, the two resulting triplets, which incorporate the concepts “artist” and “person”, semantically align well with relevant triplets incorporating abstracted entities in the KG.
Specifically, given a question $q$, we utilize an LLM to generate a structured abstract concept path $P_{q}$. The overall process is expressed as follows.
$Rat,P_{q}=\text{LLM}(q,Inst_{sa},E_{sa}),$ (4)
where $Inst_{sa}$ describes the abstraction task in a chain-of-thought (CoT) manner (Wei et al., 2023), $E_{sa}$ denotes multiple exemplars. $Rat$ denotes the rationales that represent the thought process and underpin the generation of $P_{q}$, where both are concurrently produced from a single process.
3.5 Abstraction-driven Retrieval Module
The Abstraction-driven Retrieval module retrieves question-relevant triplets as the evidence to support the answer generation. It helps reduce noise data while decreasing the context length input into the LLM.
Given a question $q$, $n$ initial topic entities $E_{q}$ and the knowledge graph $KG$, the retrieval phase over the KG is organized into multiple iterative cycles (retrieval iterations). We define two retrieval parameters: width $W$ and depth $D$. Width $W$ denotes the maximum number of topic entities and retrieved relations at each iteration, and depth $D$ represents the maximum number of iterations. Each iteration explores the 1-hop neighbor subgraphs of the topic entities and updates them with the discovered neighbors.
After retrieving $n\times W$ relevant entity-relation pairs through relation retrieval, previous RAG-based methods use LLMs to retrieve additional entities but perform poorly in the privacy-protected scenario. Instead, we first construct candidate triplets in the Relation-centric Abstraction module. Then, we retrieve question-relevant triplets and identify newly discovered entities within these triplets based on their semantic similarity to the triplets in the aforementioned abstract concept paths.
Specifically, as shown in Part (c) of Figure 2, after the Relation-centric Abstraction module, we gather a set of candidate triplets $T_{q}^{abs}$. Then, we retrieve $W$ most relevant triplets $T_{q,\text{opt}}$ from this set, based on their cosine similarity with the triplets $t_{p}$ in the abstract concept path $P_{q}$. The procedure is outlined as follows.
$T_{q,\text{opt}}^{abs}=\mathop{\arg\max}_{\begin{subarray}{c}T^{abs}_{q,\text{opt}}\subseteq T_{q}^{abs}\\ |T^{abs}_{q,\text{opt}}|\leq W\end{subarray}}\sum_{\begin{subarray}{c}t^{abs}\in T^{abs}_{q,\text{opt}}\\ t_{p}\in P_{q}\end{subarray}}\cos(\text{Emb}(t^{abs}),\text{Emb}(t_{p})).$ (5)
In each retrieval iteration, we retrieve question-relevant triplets $T^{abs}_{q,opt}$ and aggregate all triplets to date into a cumulative set $T^{abs}_{q,all}$. This set is then fed into the Generator module. Additionally, the topic entities are replaced with the newly identified entities from $T^{abs}_{q,opt}$ for the next retrieval iteration.
3.6 Generator Module
The Generator module infers the answers by feeding the question and retrieved relevant triplets into the LLM-based generator, which is similar to that of Sun et al. (2024).
Specifically, as shown in Part (d) of Figure 2, after the Abstraction-driven Retrieval module, we gather a set of question-relevant triplets $T^{abs}_{q,all}$ as evidence. Next, we input the evidence, along with the original question $q$, the instruction $Inst_{g}$ and several exemplars $E_{g}$ into an LLM-based generator. The generator then produces a flag $Flag$ and the final answers $Ans$. The process is illustrated as follows.
$Flag,Ans=\text{LLM}(T^{abs}_{q,all},q,Inst_{g},E_{g}).$ (6)
We output the answers directly if $Flag$ is positive; otherwise, we proceed to the next retrieval iteration. These answers may contain entities anonymized as MIDs. These MIDs are replaced with their real names on the user side, which saves a privacy map linking MIDs to their names.
4 Experiment
4.1 Experimental Setup
Datasets We conduct experiments on three KGQA datasets including WebQSP, CWQ and GrailQA. The details of the datasets’ splits and the number of selected samples are provided in Appendix A.1. WebQSP (Yih et al., 2016) contains questions from WebQuestions (Berant et al., 2013) that are answerable by Freebase (Bollacker et al., 2008). CWQ (Talmor and Berant, 2018) extends WebQSP, where most questions require at least 2-hop reasoning. GrailQA (Gu et al., 2021) is a diverse KGQA dataset, where most questions necessitate long-tail knowledge.
Experimental Settings We define two experimental settings. The first is the #Total setting, where we conduct the test on our complete sample set. The second is the #Filtered setting, which simulates a strict privacy-protected scenario. In this scenario, answering questions requires access to private knowledge not included in the LLM’s pre-trained data. To model this scenario, we remove all samples that can be answered correctly via CoT prompting and conduct the test on the remaining filtered subset.
Metrics We use exact match accuracy (Hits@1) as evaluation metric, consistent with previous studies (Li et al., 2023; Sun et al., 2024; Chen et al., 2024).
Implementation Details We employ gpt-4o-mini-2024-07-18 as the underlying LLM. For the Structure-oriented Abstraction module and the Generator model, the temperature parameter is set to 0 to ensure deterministic outputs. For the Relation-centric Abstraction module, the temperature is set to 0.4 to introduce variability. Both the frequency penalty and presence penalty are set to 0. The width $W$ and depth $D$ of the retrieval are set to 3, striking a balance between performance and efficiency. We repeat the experiment 3 times and report the average scores. More implementation details are illustrated in the Appendix A.3.
4.2 Baselines
Methods for applying LLMs to the KGQA task are generally divided into three categories by existing studies: Pure-LLM, SP-based and RAG-based methods. To ensure a comprehensive evaluation, we select SoTA methods from each category as our baselines, which are listed as follows.
Pure-LLM methods We select IO (Brown et al., 2020), CoT (Wei et al., 2023) and CoT-SC (Wang et al., 2023) as reference. Because they operate without external KGs and do not tailor for KGQA, they serve as baselines for evaluating the intrinsic reasoning capabilities of LLMs.
SP-based methods We select KB-BINDER (Li et al., 2023) and TrustUQA (Zhang et al., 2025) as representative SP-based methods. For their variants with dynamically exemplars, we claim them as KB-BINDER-R and TrustUQA-R, respectively.
RAG-based methods We select ToG (Sun et al., 2024), PoG (Chen et al., 2024) and GoG (Xu et al., 2024) as representative RAG-based methods. ToG (Sun et al., 2024) uses LLMs to retrieve knowledge triplets via a beam search approach. PoG (Chen et al., 2024) enhances the retrieval process through reflection, memory, and adaptive breadth. In a different vein, GoG (Xu et al., 2024) leverages LLMs as a flexible knowledge source, complementing the KG.
4.3 Performance Comparison
Type Method WebQSP CWQ GrailQA
#Tot #Fil #Tot #Fil #Tot #Fil
Pure IO 68.6 13.8 50.9 10.9 31.5 5.8
CoT 67.2 0.0 55.1 0.0 35.5 0.0
CoT-SC 68.3 8.9 54.2 2.0 35.0 3.3
SP KB-BINDER 15.8 16.0 - - 45.0 45.3
KB-BINDER-R 37.0 11.7 - - 62.3 62.8
TrustUQA 47.4 48.4 - - - -
TrustUQA-R 60.9 53.1 - - - -
RAG ToG 64.9 8.2 54.1 4.9 38.9 17.0
GoG 62.3 26.9 48.1 16.5 29.1 12.4
PoG 61.4 12.6 49.8 16.3 49.7 36.1
ARoG 74.7 58.9 60.0 36.3 78.7 71.8
Imp. $\uparrow$6.1 $\uparrow$5.8 $\uparrow$4.9 $\uparrow$19.8 $\uparrow$16.4 $\uparrow$9.0
Table 1: Comparison results (%) of Pure-LLM methods (Pure), SP-based methods (SP), and RAG-based methods (RAG), under the #Total setting (#Tot) and #Filtered setting (#Fil). The “-” indicates that the SP-based method cannot be run due to lacking special formal query annotations. “Imp.”: the absolute improvement of ARoG relative to the second-best performer. Bold: the best score. Underline: the second best score.
We compare ARoG with the baselines to demonstrate the effectiveness for question answering, the results are represented in Table 1. Overall, ARoG achieves the best performance across all three datasets under both settings.
First, ARoG significantly outperforms all RAG-based baselines. Existing RAG methods struggle with KGs containing anonymous entities, as they rely heavily on specific entity information for effective retrieval and reasoning.
Second, ARoG also surpasses existing SP-based baselines. This is because ARoG leverages LLMs to directly participate in the reasoning process through iterative retrieval-and-generation mechanisms. In contrast, SP-based baselines lack such iterative paradigms and are constrained by the coverage and completeness of the KG.
Third, compared to Pure-LLM baselines, most RAG-based baselines show no advantage, primarily due to their inability to effectively retrieve entity information in the privacy-protected scenario. In contrast, ARoG demonstrates significant improvements by effectively retrieve question-relevant triplets from the KG.
Finally, from the #Total setting to the #Filtered setting, the internal knowledge of LLMs is not enough to answer questions. Therefore, the performance of most RAG-based methods drops significantly, since they fail to retrieve private knowledge from the KG. However, ARoG compensates for this limitation with two abstraction strategies, which helps the LLMs perform reasoning over KGs without concrete entity information. For SP-based methods, since they do not require concrete entity information from the KG, their performance remains relatively stable as well.
We also conduct experiments on Qwen3-32B-FP8. The experiments yield findings that are consistent with the aforementioned conclusions. The details are shown in Appendix B.
4.4 Ablation Study
To assess the effectiveness of the two abstraction strategies in ARoG, we perform ablation studies on three datasets. Specifically, ablation studies are conducted separately on the three steps of the Relation-centric Abstraction module and on the Structure-oriented Abstraction module. Notably, relation retrieval (Step 1) aims to retrieve relations and serves as the foundation for all subsequent steps; relation filtering (Step 2) is a prerequisite for entity abstraction (Step 3). Therefore, in the ablation studies, we do not consider the scenarios of removing Step 1 or removing Step 3 while retaining Step 2. The results are shown in Table 2.
ARoG KGQA
RA SA WebQSP CWQ GrailQA
Step1 Step2 Step3 SA #Tot #Fil #Tot #Fil #Tot #Fil
$\checkmark$ $\times$ $\times$ $\times$ 63.9 53.6 36.4 20.7 72.6 66.9
$\checkmark$ $\times$ $\times$ $\checkmark$ 71.6 52.7 54.5 30.1 75.2 68.0
$\checkmark$ $\times$ $\checkmark$ $\times$ 68.3 57.4 47.9 26.9 76.9 70.7
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ 72.9 58.9 59.5 35.4 78.3 71.6
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 68.1 56.8 50.3 30.5 76.3 70.7
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.7 58.9 60.0 36.3 78.7 71.8
Table 2: Results (%) of the ablation study. RA: Relation-centric Abstraction module. SA: Structure-oriented Abstraction module. #Tot: the #Total setting. #Fil: #Filtered setting. Bold: the best score.
Analysis on Relation-centric Abstraction Module
When omitting the Relation-centric Abstraction module, we observe a performance decline of at least 2.5% under the #Total setting and at least 3.8% under the #Filtered setting. The decline is most significant under the #Filtered setting, where answering questions requires acquiring private knowledge not found in the LLM’s pre-trained knowledge. This suggests that the Relation-centric Abstraction module is highly effective at extracting and abstracting in-context semantic information from relations within KG, thereby preserving the private knowledge.
Analysis on Structure-oriented Abstraction Module
When omitting the Structure-oriented Abstraction module, we observe a performance decline of at least 2.4% under the #Total setting and at least 1.1% under the #Filtered setting. The decline is most significant on the CWQ dataset, whose questions primarily require multi-hop reasoning. This suggests that the Structure-oriented Abstraction module excels at analyzing the structural semantics of questions, while the abstract concept paths enhance multi-iteration retrieval.
4.5 Efficiency Study
We study the efficiency of RAG-based methods. Table 3 reports the average LLM calls and token consumption by these methods to answer a question across three datasets under the #Total setting.
Dataset Method #Call #Input #Output #Total
WebQSP ToG 15.4 6,356.4 1,357.0 7,713.4
GoG 10.6 12,031.0 343.3 12,374.3
PoG 10.1 5,872.5 336.0 6,208.6
ARoG 17.3 7,044.7 707.3 7,752.1
CWQ ToG 21.7 8,372.3 1,852.2 10,224.5
GoG 13.5 14,657.0 463.8 15,120.8
PoG 16.9 8,730.6 423.2 9,153.7
ARoG 25.4 1,0187.0 1,055.0 11,242.1
GrailQA ToG 13.5 5,220.1 1,306.5 6,526.5
GoG 11.4 12,665.7 428.0 13,093.7
PoG 15.8 7,445.0 436.4 7,881.4
ARoG 12.5 5,044.7 560.7 5,605.5
Table 3: Efficiency comparison between RAG-based methods under the #Total setting. #Call denotes the number of LLM calls, while #Input, #Output, #Total represent the number of input tokens, output tokens and total tokens, respectively.
On the WebQSP and CWQ datasets, we observe that ARoG is slightly less efficient than ToG and PoG. However, this disadvantage can be disregarded to some degree, especially considering ARoG’s outstanding performance. In contrast, the situation is different on the GrailQA dataset, whose most questions rely on external knowledge to answer. The reason is that ARoG can effectively utilize external knowledge while other RAG-based methods cannot, leading to higher consumption during the retrieval phase.
Additionally, we observe that GoG needs fewer LLM calls than most of others, but its input tokens consumption is too high. This is primarily due to its complex prompt design, which aims to decide whether to search knowledge from KGs or use LLMs to generate knowledge.
4.6 Quantitative Study
To understand the impact of retrieval depth $D$ and width $W$ on ARoG’s performance, we conduct the quantitative study. We vary these parameters from 1 to 4 (with a default of 3) across three datasets, following the approach in ToG (Sun et al., 2024). The results are presented in Figure 3.
Figure 3: Performance (%) of ARoG with different retrieval widths and depths.
We observe that, on the WebQSP dataset, the width $W$ is particularly critical, while the depth $D$ has little to no impact, as most answer entities can be found within the 1-hop subgraph of initial topic entities. Conversely, for the CWQ and GrailQA datasets, both width and depth enhance performance; however, the benefits diminish beyond a depth of 2, since most answers are located within the 2-hop subgraph.
4.7 Deep Analysis
Comparison between Different Abstraction Strategies
To thoroughly validate the effectiveness of abstraction strategy in the Structure-oriented Abstraction module, we compare ARoG against two alternative abstraction strategies as follows. (1) Chain-of-Thought (CoT) (Wei et al., 2023), which replaces the abstract concept path with generated rationales, and (2) question decomposition (Dec) (Chen et al., 2024), which uses derived sub-questions. For a more comprehensive baseline, we also include the variant that omits the Structure-oriented Abstraction module (“w/o SA”). The performance of all these methods is illustrated in Figure 4.
Figure 4: Performance (%) with different abstraction strategies for the Structure-oriented Abstraction module.
We observe that ARoG demonstrates superior performance across all conditions. The CoT succeeds by providing necessary entities, while the Dec succeeds by leveraging structured sub-questions. However, both strategies fail to achieve optimal results since they do not incorporate abstract concepts, which are crucial for the retrieval.
Performance in Different Privacy-Protected Scenarios
To further showcase ARoG’s robustness, we establish four different privacy scenarios: full privacy-protected scenarios (Private), exposing grounded entities information in retrieval phase (P-R), in generation phase (P-G), or both phases (P-RAG). We select ToG to represent previous RAG-based methods. We develop ARoG-R that uses the LLMs to retrieve question-relevant entities, thereby replacing the Abstraction-driven Retrieval module in ARoG. The experimental results are illustrated in Figure 5.
Figure 5: Performance (%) of ARoG, ARoG-R and ToG in different privacy-protected scenarios.
We observe that, when entities are anonymous in the generation phase, ToG’s performance drops significantly while ARoG and ARoG-R remain. The reason is that the anonymous entities retrieved by ToG are meaningless, while ARoG and ARoG-R address this issue based on the relation-centric abstraction. Additionally, when entities are anonymous in the retrieval phase, ARoG-R drops significantly while ARoG remains high. Although ToG also maintains its performance, it shows a relatively poor result. The reason is that ToG and ARoG-R retrieve the anonymous entities by feeding the question along with anonymous entities to LLMs. This approach results in suboptimal performance. In contrast, ARoG retrieve the question-relevant anonymous entities based on the Abstraction-driven Retrieval module.
Overall, ARoG exhibits impressive robustness and high performance across all three datasets and scenarios.
Case Study
The detailed case study is shown in the Appendix C.
5 Conclusion
In this paper, we investigate the privacy-protected RAG scenario in KGQA for the first time, where the entities within KGs are anonymous to LLMs. To address the challenges in this scenario, we propose a novel ARoG framework that incorporates relation-centric and structure-oriented abstraction strategies. Based on these two strategies, ARoG is able to effectively retrieve question-relevant knowledge triplets while preventing data privacy. Extensive experiments conducted on three datasets demonstrate the effectiveness and robustness of ARoG, highlighting its potential to address the challenging privacy-protected RAG scenario for KGQA.
Appendix
Appendix A Experiment Setup
A.1 Datasets
Table 4 details the dataset splits and the number of samples selected for our experiments. A notable distinction applies to the CWQ dataset, where the original annotations provide only single answer entity for each question. To facilitate a fair evaluation, we expand the answer set by executing the provided SPARQL (SPARQL Protocol and RDF Query Language) queries to retrieve all grounded entities. This methodology differs from that of ToG, GoG, and PoG, leading to a discrepancy in the reported performance on the CWQ dataset.
To ensure a fair comparison, both our proposed method and all baseline models are evaluated on the same knowledge graph. Consistent with the ToG framework, we utilize the Freebase public data dump (i.e. https://developers.google.com/freebase?hl=en), which we deploy using the Virtuoso RDF database system.
Dataset #(Train/Dev/Test) #Total #Filtered
WebQSP 3,098/-/1,639 1,639 535
CWQ 27,639/3,519/3,531 1,000 449
GrailQA 44,337/6,763/13,231 1,000 645
Table 4: Statistics of the datasets.
A.2 Settings for Baselines
IO, CoT and CoT-SC
For a fair comparison, we use the reproduced code for the IO, CoT, and CoT-SC methods from the ToG repository to infer answers.
KB-BINDER
We use the official repository and use KB-BINDER (6) (with majority vote) to infer answers.
TrustUQA
While we use the official repository for inference, we reconstruct the Condition Graph using the same knowledge graph as our method and other baselines to ensure a fair comparison. It is important to note that our reconstructed graph differs from the original in two significant ways: it is considerably larger in scale and exhibits a higher degree of noise.
ToG, GoG and PoG
For the baseline methods ToG, GoG, and PoG, we infer answers using their official repositories, adhering strictly to the default settings provided. To ensure a fair and consistent comparison, we evaluate all methods, including our proposed ARoG, using the evaluation protocol established in the official repository of ToG. Additionally, to investigate performance under a privacy-protected RAG scenario, we modify the Entity Name Search step, which prevents the leakage of sensitive entity information.
A.3 Implementation Details
In our method and all baselines, we employ gpt-4o-mini-2024-07-18 as the underlying LLM for a fair comparison, accessed via the OpenAI official API. For the Structure-oriented Abstraction module and the Generator model, the temperature parameter is set to 0 to ensure deterministic outputs. In contrast, for the Relation-centric Abstraction module, the temperature is set to 0.4 to introduce controlled variability. Both the frequency penalty and presence penalty are set to 0. The maximum token length for generation is capped at 256. Across all experiments, the width and depth of exploration are both set to 3, striking a balance between performance and efficiency. We repeat the experiment 3 times and report the average scores. All experiments are conducted on a server equipped with two Intel(R) Xeon(R) Silver 4310 CPU @ 2.10GHz and 128 GB RAM memory.
Appendix B Generalization Assessment
To validate the Generalization across different LLMs, we perform experiments using Qwen3-32B-FP8 to replace gpt-4o-mini-2024-07-18 for our method and all baselines. For a comprehensive evaluation, we benchmark ARoG against Pure-LLM and RAG-based baselines on the KGQA task. As detailed in Table 5, our method consistently outperforms these baselines.
The experiments yield findings that are consistent with the conclusions in the Section 4.3 Performance Comparison of main paper.
Type Method WebQSP CWQ GrailQA
#Tot #Fil #Tot #Fil #Tot #Fil
Pure IO 57.6 19.2 43.9 14.2 26.9 8.1
CoT 57.8 0.0 46.3 0.0 28.7 0.0
SC 57.7 3.3 46.1 1.3 29.1 2.4
SP KB-BINDER 20.2 9.1 - - 38.5 39.1
KB-BINDER-R 44.9 39.3 - - 60.5 60.9
TrustUQA 49.2 52.5 - - - -
TrustUQA-R 54.1 59.2 - - - -
RAG ToG 57.6 22.1 49.5 15.4 37.4 18.4
GoG 54.7 26.1 40.2 15.1 27.9 14.0
PoG 39.5 22.3 32.3 9.6 51.8 41.8
ARoG 72.7 64.3 54.3 36.5 78.7 74.8
Imp. $\uparrow$14.9 $\uparrow$5.1 $\uparrow$4.8 $\uparrow$21.1 $\uparrow$18.2 $\uparrow$13.9
Table 5: Comparison results (%) of Pure-LLM methods (Pure), SP-based methods (SP), and RAG-based methods (RAG), under the #Total setting (#Tot) and #Filtered setting (#Fil), using QWEN3-32B-FP8. The “-” indicates that the SP-based method cannot be run due to lacking special formal query annotations. “Imp.”: the absolute improvement of ARoG relative to the second-best performer. Bold: the best score. Underline: the second best score.
Appendix C Case Study
Figure 6 presents a typical case from the testing results on the CWQ dataset. We compare the results of ARoG, CoT, ToG, GoG and PoG in answering the question “What does the artist that was nominated for ‘The Audacity of Hope’ have a degree in?”. The underlying LLMs used are all GPT-4o-mini.
The case study illustrates the distinct operational mechanisms of each approach. In the first iteration of ARoG, it retrieves a set of question-relevant relations, such as “book.written_work.author”. It then identifies the corresponding entity cluster, which includes “m.02mjmr”. For this entity, ARoG searches and filters its adjacent relations, and then utilize an LLM to abstract it into “m.02mjmr (person)”. Concurrently, the Structure-oriented Abstraction module abstracts the question into the abstract concept path: “Common (artist) $\rightarrow$ nominated for $\rightarrow$ The Audacity of Hope (work); Common (artist) $\rightarrow$ has degree in $\rightarrow$ Communications (field of study)”. After that, the Abstraction-driven Retrieval module retrieves the question-relevant triplets, such as “The Audacity of Hope, book.written_work.author, m.02mjmr (person)”. All of the retrieved triplets are provided to the Generator module, and it generates a negative flag. Then, the topic entity is updated to “m.02mjmr (person)”. Starting from this new topic entity, the Abstraction-driven Retrieval module successfully retrieves “m.02mjmr (person), people.person.education-education.education.major_field_of_study, m.062z7 (academic discipline)”. Finally, the Generator module infers the correct answer “m.062z7 (academic discipline)”, i.e. Political Science.
In contrast, CoT relies on the LLM’s internal knowledge, which is incorrect in this case, it produces a wrong answer. Besides, ToG fails to identify the relevant anonymous entity “m.062z7” and instead retrieves incorrect triplets “m.02mjmr, award.award_nominee.award_nominations-award.award_nomination.award, m.0nh4p7s”. This failure occurs because ToG can not distinguish the entity “m.062z7” from other entities, such as “m.0nh4p7s”. Ultimately, ToG can not infer an answer and defaults to the result from the CoT approach. Similarly, although GoG and PoG introduce LLMs’ internal knowledge to supplement or modify the retrieved triplets, they introduce errors.
From this analysis, it is evident that ARoG outperforms CoT, ToG, GoG and PoG. ARoG success in retrieving question-relevant triplets and infer the correct answers, utilizing two abstraction strategies.
Figure 6: A typical case to compare different methods to answer the complex question under the privacy-protected RAG scenario for KGQA. Incorrect entities and wrong answers are highlighted in red. The noun concept of entities in the abstract concept path and retrieved triplets are highlighted in blue (question-relevant entities) or green (answer entities).
Appendix D Prompts
Here, we provide all the prompts used in ARoG. The specific in-context few-shot is shown in code files.
D.1 Relation-centric Abstraction
% The prompt used in the relation retrieval.
Please retrieve %s relations that contribute to the question from the following relations (separated by semicolons).
Q: Name the president of the country whose main spoken language was Brahui in 1980?
Topic Entity: Brahui Language
Relations: language.human_language.main_country; language.human_language.language_family; language.human_language.iso_639_3_code; base.rosetta.languoid.parent; language.human_language.writing_system; base.rosetta.languoid.languoid_class; language.human_language.countries_spoken_in; kg.object_profile.prominent_type; base.rosetta.languoid.document; base.ontologies.ontology_instance.equivalent_instances; base.rosetta.languoid.local_name; language.human_language.region
The output is:
[’language.human_language.main_country’,’language.human_language.countries_spoken_in’,’base.rosetta.languoid.parent’]
Q: {}
Topic Entity: {}
Relations: {}
% The prompt used in the entity abstraction. The ‘‘{{OBJECT_RELATIONS}}’’ and ‘‘{{SUBJECT_RELATIONS}}’’ are replaced by the related relations generated by the relation filtering.
Please provide the type with a short description of the ENTITY based on the context.
Here is an example:
Context:
ENTITY is the object of verbs: City/Town
ENTITY is the subject of verbs: Newspapers; Time zone(s); Area; Contained by; Population
The output is: {"type": "geographic location", "description": "a geographic location or administrative area that is associated with specific characteristics, such as a newspaper circulation areas, time zones, and population statistics."}
Context:
ENTITY is the object of verbs: {{OBJECT_RELATIONS}}
ENTITY is the subject of verbs: {{SUBJECT_RELATIONS}}
Now you need to directly output its type with a short description. Fill in the format below without other information or notes:
{"type": "the type of the entity", "description": "the short description of the entity"}
The output is:
D.2 Structure-oriented Abstraction
Please answer the questions step by step, the format of the reasoning process should be kept consistent with examples.
Question: What state is home to the university that is represented in sports by George Washington Colonials mens basketball?
Thought: First, the education institution has a sports team named George Washington Colonials mens basketball in is George Washington University. Second, George Washington University is in Washington D.C.. The answer is Washington, D.C..
Reasoning Path: {George Washington University (education institution) -> represented in sports by -> George Washington Colonials mens basketball (sport team)}; {George Washington University (education institution) -> located in -> Washington D.C. (state)}
Answer: Washington D.C. (state)
other in-context few-shot...
Question: {}
D.3 Generator
% The prompt used to output the flag.
Given a question and the associated retrieved knowledge graph triplets (subject entity, relation, object entity), you are asked to answer whether it is sufficient for you to answer the question with these triplets (Yes or No). Note that while several entities are represented as MIDs and corresponding type (e.g., Taylor Swift could be represented by m.0011 (person)’), you should assume the formal names are known.
Q: Find the person who said "Taste cannot be controlled by law", what did this person die from?
Knowledge Triplets: Taste cannot be controlled by law., media_common.quotation.author, m.0wfjf99 (person)
A: {No}. The retrieved knowledge graph triplet provides a quotation and identifies the author as a person (m.0wfjf99), but it does not provide any information about the cause of death of that person. To answer the question about what the person died from, additional information regarding the individuals death is required, which is not present in the provided triplet. Therefore, the triplet is insufficient to answer the question.
other in-context few-shot...
Q: {}
% The prompt used to output the answers.
Given a question and the associated retrieved knowledge graph triplets (entity, relation, entity), you are asked to answer the question with these triplets. Please note that while other entities are represented as IDs (e.g., m.0011’), you should assume you are familiar with their formal names.
Q: The artist nominated for The Long Winter lived where?
Knowledge Triplets: The Long Winter, book.written_work.author, m.0bvl_7 (person)
m.0bvl_7 (person), people.person.places_lived|people.place_lived.location, m.0wfjc51 (place)
A: {m.0wfjc51 (place)}. The artist nominated for The Long Winter, m.0bvl_7 (person), lived in a location that is represented by the entity m.0wfjc51.
other in-context few-shot...
Q: {}
Appendix E Search SPARQL
To automatically process the KG data in ARoG, we predefine the SPARQL for Freebase queries, which can be executed by filling in the entity’s mid and relation. Notably, we flatten Compound Value Type (CVT) nodes in Freebase into multiple single-hop relations in order to narrow the gap between candidate triplets and abstract concept paths.
E.1 Relations Search
% search all adjacent relations of a specific head entity.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?relation
WHERE {
ns:%s ?relation ?x .
% search all adjacent relations of a specific tail entity.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?relation
WHERE {
?x ?relation ns:%s .
E.2 Entity Cluster Search
% search all adjacent entity cluster of a specific head entity-relation pair.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
ns:%s ns:%s ?tailEntity .
% search all adjacent entity cluster of a specific tail entity-relation pair.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
?tailEntity ns:%s ns:%s .
% search all adjacent entity cluster of a specific head entity-relation pair, where the relation is flattened.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
ns:%s ns:%s ?mid_entity .
?mid_entity ns:%s ?tailEntity .
% search all adjacent entity cluster of a specific tail entity-relation pair, where the relation is flattened.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
?tailEntity ns:%s ?mid_entity .
?mid_entity ns:%s ns:%s .
E.3 Entity Name Search
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT DISTINCT ?tailEntity
WHERE {
FILTER (!isLiteral(?tailEntity) OR lang(?tailEntity) = ’’ OR langMatches(lang(?tailEntity), en’))
?entity ns:type.object.name ?tailEntity . FILTER(?entity = ns:%s)